OpenStudy (anonymous):

what is the value of (63+9)/3^2

OpenStudy (anonymous):

72/9 = 8

OpenStudy (anonymous):

8

OpenStudy (anonymous):

using exponents rule we can also solve this

OpenStudy (anonymous):

72 = 9*8 = 3^2 * 2^3 (3^2*2^3)/(3^2) = 2^3=8